# Inference-time Scaling of Diffusion Models through Classical Search

### v1

@@ C.5 Implementing Local Search with TFG hyper-parameter space
Due to the equivalence between annealed Langevin MCMC and training-free guidance with recurrence, we can implement local search with Langevin MCMC using the TFG framework of [94], efficiently searching the hyperparameters. Here we provide a overview of the algorithm and design space. Following Sec. C.3, every iteration of recurrence in Line. 5 is equivalent to an annealed Langevin MCMC step, thus $N_{\text{recur}}$ is equal to the number of local search steps.
@@
8:        $\bm{\Delta}_{\text{mean}}=\bm{\Delta}_{\text{mean}}+\mu_{t}\alpha_{t}\nabla_{{\bm{x}}_{0|t}}\log\tilde{f}({\bm{x}}_{0|t}+\bm{\Delta}_{\text{mean}})$trianglerightIterate $N_{\text{iter}}$ times starting from $\bm{\Delta}_{\text{mean}}=\bm{0}$
9:        ${\bm{x}}_{t-1}=\text{Sample}({\bm{x}}_{t},{\bm{x}}_{0|t},t)+\frac{\alpha_{t-1}}{\alpha_{t}}\left(\bm{\Delta}_{\text{var}}+\bm{\Delta}_{\text{mean}}\right)$triangleright Sample follows DDIM or DDPM
10:        ${\bm{x}}_{t}\sim\mathcal{N}\left(\frac{\alpha_{t}}{\alpha_{t-1}}{\bm{x}}_{t-1},\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)\bm{I}\right)$triangleright Recurrent strategy
@@
For time varying schedules $\rho_{t}$, $\mu_{t}$, we follow [94] and propose to use either the ‘increase’ schedule:
@@ D.1 BFS-Based Search
We present the pseudo code for BFS in Alg. 3. The core hyperparameters of BFS are the temperature $\tau_{t}$ and the evaluation time steps $\mathcal{S}$. These parameters enable a smooth transition from Best-of-N to best-first search. In general, a higher $\tau_{t}$ and evaluation at earlier time steps bias the process toward high-score samples, while a lower temperature and evaluation at later (smaller) time steps approximate the Best-of-N approach.
In practice, we typically evaluate at fractions of the sampling steps, with $\mathcal{S}=\left\{\frac{3}{4}T,\frac{1}{2}T,\frac{1}{4}T\right\}$. We perform a grid search over the temperature values to determine an appropriate magnitude using a small number of particles, and then apply this configuration when scaling up the number of particles. We use the ‘increase’ schedule in Eq. 13 for $\tau_{t}$, so that $\tau_{T}\leq\tau_{T-1}\leq\cdots\leq\tau_{0}$, reflecting increased confidence in the verifier evaluation at smaller time steps.
@@
  BFS input: Set of evaluation time steps $\mathcal{S}$. Temperature schedule $\tau_{t}$. Budget of particles $K$.
  Init: Random sample $K$ particles ${\bm{x}}_{T}^{k}\sim\mathcal{N}(\bm{0},\bm{I}),k=1,2,\cdots K$. Set the number of particles $K_{T}=K$
@@
     if $t\in\mathcal{S}$ then
        for $k=1,2,\cdots,K_{t}$ do
           Estimate the conditional mean: ${\bm{x}}_{0|t}^{k}=\frac{{\bm{x}}_{t}^{k}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{k},t)}{\alpha_{t}}$. Compute the verifier score $f({\bm{x}}_{0|t}^{k})$.
           Compute the number of children $n_{t}^{k}$:
           if Resample then
              $n_{t}^{k}=\texttt{Round}\left(K_{t}\frac{f({\bm{x}}_{0|t}^{k})^{\tau_{t}}}{\sum_{i=1}^{K_{t}}f({\bm{x}}_{0|t}^{i})^{\tau_{t}}}\right)$
           else if Prune then
              $n_{t}^{k}=\min\left(1,\texttt{Round}\left(K_{t}\frac{f({\bm{x}}_{0|t}^{k})^{\tau_{t}}}{\sum_{i=1}^{K_{t}}f({\bm{x}}_{0|t}^{i})^{\tau_{t}}}\right)\right)$
           end if
@@
     else
        $n_{t}^{k}=1$
     end if
     for $k=1,\cdots,K_{t}$ do
        Sample $n_{t}^{k}$ particles from ${\bm{x}}_{t}^{k}$: ${\bm{x}}_{t-1}^{j}\sim\tilde{p}_{\theta}(\cdot|{\bm{x}}_{t}^{k})\,,j=1,2,\cdots,n_{t}^{k}$
@@
     The number of particles at next noise level: $K_{t-1}=\sum_{k=1}^{K_{t}}n_{t}^{k}$
@@
  Return ${\bm{x}}_{0}=\operatorname*{argmax}_{k=1,\cdots,K_{0}}f({\bm{x}}_{0}^{k})$
@@ D.2 DFS-based search
In this section, we provide the details and pseudo code for DFS in Alg. 4. To better utilize previously explored sampling paths, we employ a buffer to store prior results. When no particles pass the threshold constraint, we retrieve the best sample from the buffer.
Similar to BFS, controlling the set of evaluation steps allows a trade-off between efficiency and accuracy. Evaluating at earlier time steps introduces higher uncertainty but enables earlier resampling. Additionally, adjusting the backtracking depth $\Delta_{T}$ governs the search scope: a small $\Delta_{T}$ reduces computation and favors local search, while a larger $\Delta_{T}$ enables broader exploration at the cost of increased computation.
In practice, we set the evaluation steps to $\mathcal{S}=\left\{\frac{1}{2}T,\frac{1}{4}T\right\}$ for image experiments to save compute, and to $\mathcal{S}=\left\{\frac{3}{4}T,\frac{3}{4}T-1,\cdots,1\right\}$ for PointMaze experiments. We set the recurrence depth to $T/2$ for image tasks and $T/4$ for PointMaze, corresponding to the denoised steps at which samples are first evaluated. The threshold schedule $\delta_{t}$ is also set to ‘increase’ as in Eq.13, enforcing tighter constraints for samples with lower noise.
In our experiments, we observed that when backtracking to $t_{\text{next}}=T$—thus fully restarting—the nonzero terminal SNR $\alpha_{T}/\sigma_{T}$ in many diffusion schedules [40] can lead to cumulative errors with repeated backtracking. Therefore, when backtracking to $t_{\text{next}}=T$, we initialize with fresh Gaussian noise.
@@
  DFS input Budget for total number of backtracking $B=K$, backtracking depth $\Delta_{T}$ and threshold $\left\{\delta_{t}\right\}_{t=1}^{T}$. Set of evaluation time steps $\mathcal{S}$.
@@
     if $t\in\mathcal{S}$ then
        Estimate the conditional mean and verifier score: ${\bm{x}}_{0|t}=\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\theta}({\bm{x}}_{t},t)}{\alpha_{t}}$, $f({\bm{x}}_{0|t})$.
@@
           Add the score-value pair to the buffer: $\texttt{buffer}(t).\texttt{add}\left(f({\bm{x}}_{0|t}):{\bm{x}}_{t}\right)$
@@
           if $f({\bm{x}}_{0|t})<\delta_{t}$ and budget $B=0$ then
              Pop the best sample from buffer: ${\bm{x}}_{t}\leftarrow\texttt{buffer}(t).\texttt{max}$triangleright select the best sample from past explorations
@@ Appendix E Experiment Details
E.1 Long Horizon Maze Planning
Figure 7: Additional examples for PointMaze Ultra Fig. 7(a) shows tha maze layout and a example trajectory rollout. Fig. 7(b) shows the failure case of model generated plan. Fig. 7(c) shows a successful generated plan after inference-time scaling.

### v2

@@ C.5 Implementing Local Search with TFG hyper-parameter space
Due to the equivalence between annealed Langevin MCMC and training-free guidance with recurrence, we can implement local search with Langevin MCMC using the TFG framework of Ye et al. (2024), efficiently searching the hyperparameters. Here we provide a overview of the algorithm and design space. Following Sec. C.3, every iteration of recurrence in Line. 5 is equivalent to an annealed Langevin MCMC step, thus $N_{\text{recur}}$ is equal to the number of local search steps.
@@
[…]
9:   ${\bm{x}}_{t-1}=\text{Sample}({\bm{x}}_{t},{\bm{x}}_{0|t},t)+\frac{\alpha_{t-1}}{\alpha_{t}}\left(\bm{\Delta}_{\text{var}}+\bm{\Delta}_{\text{mean}}\right)$ $\triangleright$ Sample follows DDIM or DDPM
10:   ${\bm{x}}_{t}\sim\mathcal{N}\left(\frac{\alpha_{t}}{\alpha_{t-1}}{\bm{x}}_{t-1},\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)\bm{I}\right)$ $\triangleright$ Recurrent strategy
@@
For time varying schedules $\rho_{t}$, $\mu_{t}$, we follow Ye et al. (2024) and propose to use either the ‘increase’ schedule:
@@ D.1 BFS-Based Search
We present the pseudo code for BFS in Alg. 3. In practice, we can evaluate and resample the particles at a fixed subset of time steps, with detailed hyper-parameters in the details of each experiment.
@@
BFS input: Tempering schedule $\tau_{t}$. Budget of particles NN. Scoring function and resampling method.
Init: Random sample NN particles ${\bm{x}}_{T}^{k}\sim\mathcal{N}(\bm{0},\bm{I}),k=1,2,\cdots N$.
@@
  for $k=1,2,\cdots,N$ do
Estimation
Estimate the conditional mean: ${\bm{x}}_{0|t}^{k}=\frac{{\bm{x}}_{t}^{k}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{k},t)}{\alpha_{t}}$. Compute the verifier score estimate $f({\bm{x}}_{0|t}^{k})$.
Scoring
Score the particles according to the scoring functions:
$\textbf{Current}:\widehat{f}({\bm{x}}_{t}^{k})=\tau_{t}f({\bm{x}}_{0|t}^{k}),$
$\textbf{Difference}:\widehat{f}({\bm{x}}_{t}^{k})=\tau_{t}f({\bm{x}}_{0|t}^{k})-\widehat{f}_{\text{prev}}^{k},$
$\textbf{Max}:\widehat{f}({\bm{x}}_{t}^{k})=\max\left(\tau_{t}f({\bm{x}}_{0|t}^{k}),\widehat{f}_{\text{prev}}^{k}\right).$
Resampling
Compute the weights
$(w_{t}^{1},w_{t}^{2},\cdots,w_{t}^{N})=\text{softmax}\left(\widehat{f}({\bm{x}}_{t}^{1}),\widehat{f}({\bm{x}}_{t}^{2}),\cdots\widehat{f}({\bm{x}}_{t}^{N})\right)\,$,
and resample the children
$(n_{t}^{1},n_{t}^{2}\cdots,n_{t}^{N})=\text{Resample}(N;w_{t}^{1},w_{t}^{2},\cdots,w_{t}^{N}),$
where Resample can be Multinomial resampling or SSP resampling (Alg.4).
@@
  for $k=1,\cdots,N$ do
   Sample $n_{t}^{k}$ children particles from ${\bm{x}}_{t}^{k}$: ${\bm{x}}_{t-1}^{j}\sim\tilde{p}_{\theta}(\cdot|{\bm{x}}_{t}^{k})\,,j=1,2,\cdots,n_{t}^{k}$
@@
  Update the score buffers for computing the Difference and Max score at next timestep.
$\widehat{f}_{\text{prev}}^{k}=\left\{\begin{aligned} &\widehat{f}\left({\bm{x}}_{t}^{\text{parent(k)}}\right)\quad\text{if}~~\text{scoring function}~~\text{is}~~\textbf{Max}\\ &\tau_{t}f\left({\bm{x}}_{0|t}^{\text{parent(k)}}\right)\quad\text{if}~~\text{scoring function}~~\text{is}~~\textbf{Difference}\end{aligned}\right$.
@@
Return ${\bm{x}}_{0}=\operatorname*{argmax}_{{\bm{x}}_{0}^{1},\cdots,{\bm{x}}_{0}^{N}}f({\bm{x}}_{0}^{k})$
D.1.1 Overview of design space
Tempering
When estimating the particle score via the denoised output $f({\bm{x}}_{0|t}^{k})$, for larger $t$ the estimation often has larger bias, due to the fact that $\mathbb{E}[f({\bm{x}}_{0})|{\bm{x}}_{t}]\neq f(\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}])=f({\bm{x}}_{0|t})$. Thus, we propose to re-weight the score estimates with a tempering schedule $\tau_{t}$ before scoring the particles. For a increasing schedule $\tau_{T}<\tau_{T-1}<\cdots<\tau_{1}<\tau_{0}$, the estimated scores have more influence at smaller time steps with higher weights, where the estimation has less bias. For the increasing tempering schedule, we simply adopt the design in DAS (Kim et al., 2025) with $\tau_{t}=\left((1+\gamma)^{T-t}-1\right)\tau$, with $\gamma$ searched in $\left\{0.008,0.024\right\}$. We also consider the deterministic particle selection in SVDD (Li et al., 2024) where we only retain the highest scoring particle, which can be seen as $\tau_{t}=\infty$. For the constant tempering schedule we simply set $\tau_{t}=\tau$, without any tempering on different time steps.
Scoring
Given the score estimates $f({\bm{x}}_{0|t})$, one can construct scoring rules based on the scores of the entire path $\left\{f({\bm{x}}_{0|s}^{k})\right\}_{s=T}^{t}$. In traditional sequential-monte-carlo (SMC) based methods, the score is computed as the difference between two consecutive evaluations: $\widehat{f}({\bm{x}}_{t}^{k})=\tau_{t}f({\bm{x}}_{0|t}^{k})-\tau_{t+1}f({\bm{x}}_{0|t+1}^{k})$. However, in FK-steering (Singhal et al., 2025a), they pointed out that the difference potential is unfair for sample path that saturates earlier, reaching a high reward at earlier time steps. Thus they propose the max scoring: $\widehat{f}({\bm{x}}_{t}^{k})=\max_{s\geq t}\tau_{t}f({\bm{x}}_{0|s}^{k})$. We also consider the simple baseline of current score: $\widehat{f}({\bm{x}}_{t}^{k})=\tau_{t}f({\bm{x}}_{0|t}^{k})$.
Resampling
Given the scores $\widehat{f}({\bm{x}}_{t}^{k})$, we can compute the softmax logits of particles via $\left\{w_{t}^{k}\right\}_{k=1}^{N}=\text{softmax}\left\{\widehat{f}({\bm{x}}_{t}^{k})\right\}_{k=1}^{N}$. Then, we resample the particles according to the logits, and obtain the number of children for each particle $n_{t}^{k}$. The sets $\left\{n_{t}^{k}\right\}_{k=1}^{N}$ should satisfy the following constraints:
$\sum_{k=1}^{N}n_{t}^{k}=$ $N$
$\mathbb{E}[n_{t}^{k}]=$ $Nw_{t}^{k}$
Most previous approaches simply adopt the multinomial resampling, with $\left\{n_{t}^{k}\right\}_{k=1}^{N}\sim\text{Multinomial}\left(N;\left\{w_{t}^{k}\right\}_{k=1}^{N}\right)$. However, this sampling process introduces large variance since every component is sampled independently. To reduce variance, for example, residual resampling only resamples the residuals $\tilde{n}_{t}^{k}$ from NRN-R with weights $\tilde{w}_{t}^{k}=\frac{Nw_{t}^{k}-[Nw_{t}^{k}]}{N-R}$, where $R=\sum_{k=1}^{N}[Nw_{t}^{k}]$. Then, we obtain $n_{t}^{k}=[Nw_{t}^{k}]+\tilde{n}_{t}^{k}$. This way we reduce the randomness in $n_{t}^{k}$.
The SSP resampling views the sampling process as randomized rounding on the expectations $Nw_{t}^{k}$, with pseudo code from Gerber et al. (2020) shown below in Alg. 4.
Algorithm 4 SSP resampling
Inputs: $u\in[0,1]^{{N}}$ and $(\xi_{1},\dots,\xi_{N})\in\mathbb{R}_{+}^{N}$ such that $\xi_{n}=Nw_{t}^{n}$ with $w_{t}^{n}$ being the softmax weight of the $n$-th particle, and we have $\sum_{n=1}^{N}\xi_{n}=N\in\mathbb{N}$.
Output: $\left(Y_{\text{ssp}}^{1}(u),\dots,Y_{\text{ssp}}^{N}(u)\right)\in\mathbb{N}^{N}$ such that $\sum_{n=1}^{N}Y_{\text{ssp}}^{n}(u)=\sum_{n=1}^{N}\xi_{n}$.
Initialization: $\left(Y_{\text{ssp}}^{1}(u),\dots,Y_{\text{ssp}}^{N}(u)\right)\leftarrow(\xi_{1},\dots,\xi_{N})$, $(n,m,k)\leftarrow(1,2,1)$
Iterate the following steps until $\left(Y_{\text{ssp}}^{1}(u),\dots,Y_{\text{ssp}}^{N}(u)\right)\in\mathbb{N}^{N}$:
(1) Let $\delta$ be the smallest number in $(0,1)$ such that at least one of $Y_{\text{ssp}}^{n}(u)+\delta$ or $Y_{\text{ssp}}^{m}(u)-\delta$ is an integer, and let $\epsilon$ be the smallest number in $(0,1)$ such that at least one of $Y_{\text{ssp}}^{n}(u)-\epsilon$ or $Y_{\text{ssp}}^{m}(u)+\epsilon$ is an integer.
(2) If $u_{k}\leq\epsilon/(\delta+\epsilon)$ set $(Y_{\text{ssp}}^{n}(u),Y_{\text{ssp}}^{m}(u))\leftarrow(Y_{\text{ssp}}^{n}(u)+\delta,Y_{\text{ssp}}^{m}(u)-\delta)$; otherwise set $(Y_{\text{ssp}}^{n}(u),Y_{\text{ssp}}^{m}(u))\leftarrow(Y_{\text{ssp}}^{n}(u)-\epsilon,Y_{\text{ssp}}^{m}(u)+\epsilon)$.
(3) Update $n$ and $m$ as follows:
  1. 1.
    If $\left(Y_{\text{ssp}}^{n}(u),Y_{\text{ssp}}^{m}(u)\right)\in\mathbb{N}^{2}$, $(n,m)\leftarrow(m+1,m+2)$;
  2. 2.
    If $Y_{\text{ssp}}^{n}(u)\in\mathbb{N}$ and $Y_{\text{ssp}}^{m}(u)\not\in\mathbb{N}$ set $(n,m)\leftarrow(m,m+1)$;
  3. 3.
    if $Y_{\text{ssp}}^{n}(u)\not\in\mathbb{N}$ and $Y_{\text{ssp}}^{m}(u)\in\mathbb{N}$ set $(n,m)\leftarrow(n,m+1)$.
(4) $k\leftarrow k+1$
There are more variance reduced sampling methods such as systematic resampling and stratified resampling, and we adopt the Srinivasan sampling process (SSP) resampling as our design choice. For other variance-reduced methods, the performance is similar to SSP resampling, all outperforming naive multinomial resampling.
D.1.2 Overview of prior methods
Here, we provide an overview of prior methods.
SVDD (Li et al., 2024)
In SVDD, the best sample is selected at each time step, from which $M$ children are generated. This approach can be viewed as a variant of BFS with $\tau=\infty$ and $M$ particles. Nodes are evaluated using the current score estimate $f({\bm{x}}_{0|t})$.
TreeG (Guo et al., 2025)
In TreeG, particles are ranked and the top $M$ are either selected directly or resampled based on their scores to obtain $M$ samples. From each selected particle, $K$ children are sampled, resulting in an effective tree width of $KM$. Particles are evaluated using their current score $f({\bm{x}}_{0|t})$. They also consider adding Gaussian noise to ${\bm{x}}_{0|t}$ to approximate the posterior distribution $p({\bm{x}}_{0}|{\bm{x}}_{t})$. However, the posterior distribution is intractable as it is not a Gaussian distribution, and exact Gaussian approximation requires computing the variance using the Jacobian of the score function (see Ye et al. (2024)).
DAS (Kim et al., 2025)
In DAS, the authors propose an exponentially increasing tempering schedule as the default, given by $\tau_{t}=(1+\gamma)^{T-t}-1$, and also introduce an adaptive tempering schedule. They adopt advanced SSP resampling instead of multinomial resampling, and evaluate particles based on the difference in rewards from the previous evaluation, similar to SMC methods.
FK-steering (Singhal et al., 2025a)
In FK, the authors propose several options for evaluating intermediate particles, including difference, max, and sum, with max adopted as the default. In the official implementation, multinomial resampling is used, which may lead to suboptimal performance.
@@ D.2 DFS-based search
In this section, we provide the details and pseudo code for DFS in Alg. 5. To better utilize previously explored sampling paths, we employ a buffer to store prior results. When the budget is exhausted and no more backtracking is allowed, we retrieve the best sample from the buffer.
Similar to BFS, controlling the set of evaluation steps allows a trade-off between efficiency and accuracy. Evaluating at earlier time steps introduces higher uncertainty but enables backtracking. Additionally, adjusting the backtracking depth $\Delta_{T}$ governs the search scope: a small $\Delta_{T}$ reduces computation and favors local search, while a larger $\Delta_{T}$ enables broader exploration at the cost of increased computation.
In practice, we set the evaluation steps to $\mathcal{S}=\left\{\frac{1}{2}T,\frac{1}{4}T\right\}$ for image experiments to save compute, and to $\mathcal{S}=\left\{\frac{3}{4}T,\frac{3}{4}T-1,\cdots,1\right\}$ for PointMaze experiments. We set the recurrence depth to $T/2$ for image tasks and $T/4$ for PointMaze, corresponding to the denoised steps at which samples are first evaluated. The threshold schedule $\delta_{t}$ is also set to ‘increase’ as in Eq.14 in our PointMaze experiments, enforcing tighter constraints for samples with lower noise. We use a simple constant $\delta_{t}=\delta$ schedule in our image experiments.
In our experiments, we observed that when backtracking to $t_{\text{next}}=T$—thus fully restarting—the nonzero terminal SNR $\alpha_{T}/\sigma_{T}$ in many diffusion schedules (Lin et al., 2024) can lead to cumulative errors with repeated backtracking. Therefore, when backtracking to $t_{\text{next}}=T$, we initialize with fresh Gaussian noise.
@@
DFS input: Budget for total number of backtracking $B=K$, backtracking depth $\Delta_{T}$ and threshold $\left\{\delta_{t}\right\}_{t=1}^{T}$.
@@
  Compute the conditional mean ${\bm{x}}_{0|t}=\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\theta}({\bm{x}}_{t},t)}{\alpha_{t}}$, and estimate the verifier score $f({\bm{x}}_{0|t})$.
  Add the score-sample pair to the buffer: $\texttt{buffer}(t).\texttt{add}\left(f({\bm{x}}_{0|t}):{\bm{x}}_{t}\right)$
@@
   if $B=0$ then
    Pop the best sample from buffer: ${\bm{x}}_{t}\leftarrow\texttt{buffer}(t).\texttt{max}$   $\triangleright$ select the best sample from past explorations
@@ Appendix E Experiment Details
E.1 Ablation of BFS design space
We directly adopt the official code base of FK-steering Singhal et al. (2025a) and use the samping methods provided in the code base of DAS (Kim et al., 2025). We use the ImageReward prompts as in (Singhal et al., 2025a) and report the average and standard deviation over 4 independent trials. For the temperature and resampling interval, we directly follow the implementation of FK-steering. For TreeG (Guo et al., 2025) we use a fixed branch out size of 2. The detailed hyper-parameters are below in Table. 6:
Table 6: BFS implementation parameters
temperature 10
resampling interval [20,40,80]
sampling steps 100
num seeds 4
E.2 Text-to-Image Compositional Generation with DFS
We use the SSD-1B model [^1] https://huggingface.co/segmind/SSD-1B which is distilled from SDXL, and we use the default sampling configuration with 50 steps of DDIM sampler. For DFS and BFS, we evaluate at time steps $\left\{25,35,45\right\}$ and set the backtrack depth $\Delta_{T}=25$. For BFS we additionally sweep the temperature in range $\left\{0.5,1,2,4,8\right\}$ and report the best performance.
We provide example visualizations in Fig. 7.
(a) w/o inference scaling
(b) w/ inference scaling
(c) Example from color dataset with prompt: a yellow book and a red vase
(d) w/o inference scaling
(e) w/ inference scaling
(f) Example from texture dataset with prompt: a fabric hat and a glass mirror
(g) w/o inference scaling
(h) w/ inference scaling
(i) Example from shape dataset with prompt: a round bagel and a square toaster
Figure 7: Example of inference-time scaling with CompBench Above shows the example prompts from CompBench benchmark and images before and after inference scaling
E.3 Long Horizon Maze Planning